OpenStudy (anonymous):

Find the indefinite integral:

OpenStudy (anonymous):

\[\frac{ (x+5)dx }{ (\sqrt{x}-3)^{\frac{ 3 }{ 2 }}}\]

OpenStudy (jhannybean):

Wow this thing requires something like a double substitution D:

myininaya (myininaya):

I was thinking of a doing a trig sub what about you @Jhannybean ?

OpenStudy (jhannybean):

double u-sub

myininaya (myininaya):

oh just 2 algebraic subs that would be a lot more cute

OpenStudy (jhannybean):

Or uh what do you call it, a u-sub and then like a z sub or something, something that replaces the u.

OpenStudy (jhannybean):

\(\boxed{`cute`}\) haha.

OpenStudy (jhannybean):

I'm just going to paste the wolfram partial solution here.

myininaya (myininaya):

and you want to look at the trig sub i think this should work \[\sqrt[4]{x}=\sqrt{3} \sec(\theta) \\ \sqrt{x}=3 \sec^2(\theta) \\ x=9 \sec^4( \theta) \\ dx=36\sec^3(\theta) \sec(\theta) \tan(\theta) d \theta \]

OpenStudy (jhannybean):

Ohh, that works out nicely it seems.